BRICS Notes Series, Abstracts, 2004

September 16, 2004

This document is also available as PostScript and DVI.

References

[NS-04-1] PDF.

Luca Aceto, Willem Jan Fokkink, and Irek Ulidowski, editors. Preliminary Proceedings of the Workshop on Structural Operational

Semantics, SOS ’04, (London, United Kingdom, August 30, 2004), August 2004.

vi+56. Abstract: This volume contains the preliminary proceedings of

SOS ’04 held in London, United Kingdom, August 30, 2004. The final proceedings will appear as a volume in the ENTCS series.